I am completely lost and would like the step by step play so I can understand what to do.
Subtract the polynomials (-5v^2+6v+3)-(8v^2+3)=

Answer by ankor@dixie-net.com(12706)

(-5v^2+6v+3)-(8v^2+3) =
remove the brackets,
remember, if there is a minus outside the brackets, you change all the signs inside the brackets
-5v^2 + 6v + 3 - 8v^2 - 3 =
:
group like terms
-5v^2 - 8v^2 + 6v + 3 - 3 =
:
Combine like terms
-13v^2 + 6v